find the measure of the angle,its complement,and its supplement when.....
twice the angle is 72degress less than the supplement of the angle.
i have writen the formula i thought is the outcome of this,but i believe its incorrect
2x=72-180-x
is this correct?
but i do understand that in order to fine to complement and the supplement you must use 90-x and 180-x.

Answer by ankor@dixie-net.com(12692)

find the measure of the angle,its complement,and its supplement when.....
twice the angle is 72degress less than the supplement of the angle.
i have writen the formula i thought is the outcome of this,but i believe its incorrect
2x=72-180-x
is this correct?
but i do understand that in order to fine to complement and the supplement you must use 90-x and 180-x.
:
Almost, it says " 72 less than the supplement" so it should be
2x = (180-x) - 72
2x = 180 - x - 72
2x + x = 180 - 72
3x = 108
x = 36